The tax rates are as shown. Your firm currently has taxable income of 79,800.

How much additional tax will you owe if you increase your taxable income by 22,300?

Taxable Income Tax Rate
0-50,000 15%
50,001 - 75,000 25%
75,001 - 100,000 34%
100,001 - 335,000 39%
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Previously tax =(0.15*50,000)+(0.25*(75000-50,000))+(0.34*(79800 - 75000)

=$15382

Now taxable income=(79800+22300)=$102,100

Hence tax =(0.15*50,000)+(0.25*(75000-50,000))+(0.34*(100,000-75000))+(0.39*2100)

=$23069

hence additional tax=(23069-15382)=$7687